# Representability of the direct sum of uniform $q$-matroids

### v1

@@ 1.3. The direct sum of q-matroids
In [10] the notion of direct sum of two $q$-matroids has been introduced. We are going to define the direct sum of $t$ many $q$-matroids iteratively, thanks to the associativity property of the direct sum, proved in [19]. We use Notation 1.7.
@@
Theorem 1.15 ([17, Theorem 3.4]).
Let $\mathcal{M}_{i}=(\mathbb{F}_{q}^{n_{i}},\rho_{i}),i=1,2,$ be $q$-matroids of rank $k_{i}$, and let $\mathcal{M}=\mathcal{M}_{1}\oplus\mathcal{M}_{2}$. Suppose $\mathcal{M}$ is representable over $\mathbb{F}_{q^{m}}$. Then $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are representable over $\mathbb{F}_{q^{m}}$ and $\mathcal{M}=(\mathcal{S}_{1}\oplus\mathcal{S}_{2},\rho_{\mathcal{S}_{1}\oplus\mathcal{S}_{2}})$ for some $[n_{i},k_{i}]_{q^{m}/q}$ systems $\mathcal{S}_{i}$, $i\in\{1,2\}$. Furthermore, $(\mathcal{S}_{i},\rho_{\mathcal{S}_{i}})$ is an $\mathbb{F}_{q^{m}}$-representation of $\mathcal{M}_{i}$ for each $i\in\{1,2\}$.
@@
Example 1.17 ([17, Proposition 3.7]).
Let $\mathbb{F}_{q}=\mathbb{F}_{2}$ and $\mathbb{F}_{4}=\{0,1,\omega,\omega+1\}$. Consider the full $[4,2]_{4/2}$ system $\mathcal{S}=\mathbb{F}_{4}^{2}$ and set $\mathcal{M}:=(\mathcal{S},\rho_{\mathcal{S}})$. Then $\mathcal{M}\oplus\mathcal{M}$ is not representable.
@@
Let $\mathcal{M}=(\mathbb{F}_{q}^{n},\rho)$ be a $q$-matroid of rank $k$, with $0<k<n$. Then $\mathcal{M}=\mathcal{U}_{k,n}(q)$ if and only if $\mathcal{M}$ has only two cyclic flats, namely $\langle 0\rangle$ and $\mathbb{F}_{q}^{n}$.
@@ 2. The direct sum of uniform q-matroids
In this section we take the first steps toward the geometric characterization of the representability of the direct sum of $t$ uniform $q$-matroids $\mathcal{U}_{k_{1},n_{1}}\oplus\cdots\oplus\mathcal{U}_{k_{t},n_{t}}$, by characterising its independent spaces. In the rest of the paper we will assume that none of the $t$ summands is the trivial or the free uniform $q$-matroid. This is due to the following observation.
@@
Furthermore, in [17, Proposition 3.9] it is proven that a $q$-matroid $\mathcal{M}$ is $\mathbb{F}_{q^{m}}$-representable as $\mathcal{M}=(\mathcal{S},\rho_{\mathcal{S}})$, if and only if $\mathcal{M}\oplus\mathcal{U}_{s,s}(q)$ is $\mathbb{F}_{q^{m}}$-represented by $(\mathcal{S}\oplus\mathbb{F}_{q}^{s},\rho_{\mathcal{S}\oplus\mathbb{F}_{q}^{s}})$.
@@
For the rest of the paper, assume $1\leq k_{i}<n_{i}\leq n$ and let $\mathcal{U}_{k_{i},n_{i}}(q)$, $i\in[t]$ be uniform $q$-matroids.
@@
Let $k_{1},\ldots,k_{t},n_{1},\ldots,n_{t},m,r$ be positive integers, with $1\leq k_{i}<n_{i}\leq m$ for $i\in[t]$. Let $\mathcal{M}=\mathcal{U}_{k_{1},n_{1}}(q)\oplus\cdots\oplus\mathcal{U}_{k_{t},n_{t}}(q)$ and for each $\mathcal{J}\subseteq[t]$ denote $k_{\mathcal{J}}:=\sum_{j\in\mathcal{J}}k_{j}$. Then $I\in\mathcal{I}(\mathcal{M})$ if and only if
@@
Let $\mathbf{k}=(k_{1},\ldots,k_{t})\in\mathbb{N}^{t}$ and let $h,k$ be positive integers such that $k=k_{1}+\ldots+k_{t}$ and $11\leq h\leq k-$ 1, then we define
$\Lambda_{h,\mathbf{k}}:=\left\{V:V\textnormal{ is }\mathbb{F}_{q^{m}}\textnormal{-subspace of }\mathbb{F}_{q^{m}}^{k},\dim_{\mathbb{F}_{q^{m}}}(V)=h\textnormal{ and }\iota_{i}(\mathbb{F}_{q^{m}}^{{k_{i}}})\not\subseteq V\mbox{ for }i\in[t]\right\}.$
@@
Let $k_{1},\ldots,k_{t},k,n_{1},\ldots,n_{t},m,r$ be positive integers, with $1\leq k_{i}<n_{i}\leq m$ for $i\in[t]$ and $k=k_{1}+\ldots+k_{t}$. Let $\mathcal{S}_{i}$ be a $[n_{i},k_{i}]_{q^{m}/q}$ system for $i\in[t]$. If $\mathcal{S}=\mathcal{S}_{1}\oplus\ldots\oplus\mathcal{S}_{t}$ is $(\Lambda_{k-1,\mathbf{k}},r)$-evasive, then $\mathcal{S}$ is $(\Lambda_{l,\mathbf{k}},r-k+1+l)$-evasive for any $l\in[k-1]$.
@@
We prove the statement by induction on $k-1-l$.
If $k-1-l=0$, then there is nothing to prove and the assertion is clearly true.
Suppose that $l<k-1$ and assume that the claim holds true for $l+1$. Let us take any subspace $T\in\Lambda_{l,\mathbf{k}}$. We will show that there exists a subspace $T^{\prime}\in\Lambda_{l+1,\mathbf{k}}$ containing $T$ and such that
@@
$\dim_{\mathbb{F}_{q}}(T\cap\mathcal{S})\leq\dim_{\mathbb{F}_{q}}(T^{\prime}\cap\mathcal{S})-1\leq r-k+1+l+1-1=r-k+1+l$
for every $T\in\Lambda_{l,\mathbf{k}}$.
@@
Since $\mathcal{S}_{i}$ is an $[n_{i},k_{i}]_{q^{m}/q}$ system contained in $\iota_{i}(\mathbb{F}_{q^{m}}^{{k_{i}}})$ there exists $v\in\mathcal{S}_{i}\setminus T$. Clearly, $T^{\prime}=T+\langle v\rangle_{\mathbb{F}_{q^{m}}}\in\Lambda_{l+1,\mathbf{k}}$ and satisfies (5).
Case II. Now assume that for any $i\in[t]$ we have that
@@
So, $l\geq k-t$ and $T\supseteq(T\cap\iota_{1}(\mathbb{F}_{q^{m}}^{{k_{1}}}))+(T\cap\iota_{2}(\mathbb{F}_{q^{m}}^{{k_{2}}}))+\ldots+(T\cap\iota_{t}(\mathbb{F}_{q^{m}}^{{k_{t}}}))$. Consider $T_{i}=\iota_{1}(\mathbb{F}_{q^{m}}^{{k_{1}}})+\iota_{i}(\mathbb{F}_{q^{m}}^{{k_{i}}})$ for any $i\in\{2,\ldots,t\}$. Then, $\dim_{\mathbb{F}_{q^{m}}}(T_{i})=k_{1}+k_{i}$ and
$\dim_{\mathbb{F}_{q^{m}}}(T\cap T_{i})\in\{k_{1}+k_{i}-2,k_{1}+k_{i}-1\}.$
@@
Case II.a. Assume that there exists $i\in\{2,\ldots,t\}$ such that $\dim_{\mathbb{F}_{q^{m}}}(\mathcal{S}\cap T_{i})=k_{1}+k_{i}-2$. This implies that
$T\cap T_{i}=(T\cap\iota_{1}(\mathbb{F}_{q^{m}}^{{k_{1}}}))+(T\cap\iota_{i}(\mathbb{F}_{q^{m}}^{{k_{i}}}))$
and there exist $v_{1}\in\mathcal{S}_{1}$ and $v_{i}\in\mathcal{S}_{i}$ such that $v_{1},v_{i}\notin T$. Note that $T\cap\langle v_{1},v_{i}\rangle_{\mathbb{F}_{q^{m}}}=\{0\}$ and $v_{1}+v_{i}\in\mathcal{S}_{1}\oplus\mathcal{S}_{i}$, therefore $T^{\prime}=T\oplus\langle v_{1}+v_{i}\rangle_{\mathbb{F}_{q^{m}}}\in\Lambda_{l+1,\mathbf{k}}$ and satisfies (5).
Case II.b. Suppose now that for every $i\in\{2,\ldots,t\}$ we have $\dim_{\mathbb{F}_{q^{m}}}(\mathcal{S}\cap T_{i})=k_{1}+k_{i}-1$. Since $\dim_{\mathbb{F}_{q^{m}}}(T\cap T_{i})=k_{1}+k_{i}-1$ and $\dim_{\mathbb{F}_{q}}(T\cap\iota_{i}(\mathbb{F}_{q^{m}}^{{k_{i}}}))=k_{i}-1$, then for each $i\in\{2,\ldots,t\}$ there exists $v_{i}\in T_{i}\setminus(\iota_{1}(\mathbb{F}_{q^{m}}^{{k_{1}}})\cup\iota_{i}(\mathbb{F}_{q^{m}}^{{k_{i}}}))$ such that
$T=(T\cap\iota_{1}(\mathbb{F}_{q^{m}}^{{k_{1}}}))+(T\cap\iota_{2}(\mathbb{F}_{q^{m}}^{{k_{2}}}))+\ldots+(T\cap\iota_{t}(\mathbb{F}_{q^{m}}^{{k_{t}}}))+\langle v_{2},\ldots,v_{t}\rangle_{\mathbb{F}_{q^{m}}}.$
It follows that $\dim_{\mathbb{F}_{q^{m}}}(T)=k-1$, a contradiction to the assumption that $l<k-1$. Thus, this subcase is not possible and this concludes the proof. ∎
@@
Let $k_{1},\ldots,k_{t},k,n_{1},\ldots,n_{t},n,m$ be positive integers, with $1\leq k_{i}<n_{i}\leq m$ for $i\in[t]$ and $k=k_{1}+\ldots+k_{t}$, $n=n_{1}+\ldots+n_{t}$. Let $\mathcal{S}_{i}$ be an $\mathbb{F}_{q^{m}}$-representation for $\mathcal{U}_{k_{i},n_{i}}(q)$ for each $i\in[t]$. Define the $[n,k]_{q^{m}/q}$ system
@@
$(1)\Longleftrightarrow(2)$: Let us call $\mathcal{M}_{1}=(\mathcal{S},\rho_{\mathcal{S}})$ and $\mathcal{M}_{2}=\mathcal{U}_{k_{1},n_{1}}(q)\oplus\ldots\oplus\mathcal{U}_{k_{t},n_{t}}(q)$. Then, $\mathcal{M}_{1}\cong\mathcal{M}_{2}$ if and only if there exists an invertible $\mathbb{F}_{q}$-linear map $\psi:\mathbb{F}_{q}^{n_{1}+\ldots+n_{t}}\longrightarrow\mathcal{S}$ such that $\mathcal{I}(\mathcal{M}_{1})=\mathcal{I}(\psi(\mathcal{M}_{2}))$. Let $G$ be any generator matrix associated to the $[n,k]_{q^{m}/q}$ system $\mathcal{S}$ of the form
@@
Case I. If $\Gamma$ does not contain any of the $\mathbb{F}_{q^{m}}$-subspaces $\iota_{i}(\mathbb{F}_{q^{m}}^{{k_{i}}})$, then $\ell:=\rho_{\mathcal{S}}(I)<k$ and $\Gamma\in\Lambda_{\ell,\mathbf{k}}$. This contradicts the hypothesis of $\mathcal{S}$ being $(\Lambda_{k-1,\mathbf{k}},k-1)$-evasive, due to Lemma 2.3. Furthermore, this also shows that for every $\mathbb{F}_{q}$-subspace $I^{\prime}\subseteq\mathcal{S}$, if $\Gamma^{\prime}:=\langle I^{\prime}\rangle_{\mathbb{F}_{q^{m}}}$ does not contain any space of the form $\iota_{i}(\mathbb{F}_{q^{m}}^{{k_{i}}})$ for $i\in[t]$, then
$\rho_{\mathcal{S}}(I^{\prime})=\min\{k,\dim_{\mathbb{F}_{q}}(I^{\prime})\}.$ (7)
Case II. Up to a permutation of the set $[t]$, we can assume without loss of generality that $\Gamma$ contains the $\mathbb{F}_{q^{m}}$-subspaces $\iota_{i}(\mathbb{F}_{q^{m}}^{{k_{i}}})$, for each $i\in[a]$, for some $a$ such that $0<a<t$ (note that $a=0$ is Case I, and $a=t$ was already proved before). Let $\dim_{\mathbb{F}_{q}}(I)=r+s+t$ with $r=\dim_{\mathbb{F}_{q}}(I\cap\mathcal{S}_{[a]})\leq k_{[a]}$ and $s=\dim_{\mathbb{F}_{q}}(I\cap\mathcal{S}_{[t]\setminus[a]})\leq k_{[t]\setminus[a]}$. $I$ can be written as
$I=\langle(v_{1}\mid 0),\ldots,(v_{r}\mid 0)\rangle_{\mathbb{F}_{q}}+\langle(0\mid u_{1}),\ldots,(0\mid u_{s})\rangle_{\mathbb{F}_{q}}+\langle(w_{1}\mid z_{1}),\ldots,(w_{t}\mid z_{t})\rangle_{\mathbb{F}_{q}}.$
@@
    $X_{1}:=\langle v_{1},\ldots,v_{r},w_{1},\ldots,w_{t}\rangle_{\mathbb{F}_{q}}\subseteq\mathcal{S}_{[a]}$ and $\dim_{\mathbb{F}_{q}}(X_{1})=r+t$;
@@
    $X_{2}:=\langle u_{1},\ldots,u_{s},z_{1},\ldots,z_{t}\rangle_{\mathbb{F}_{q}}\subseteq\mathcal{S}_{[t]\setminus[a]}$ and $\dim_{\mathbb{F}_{q}}(X_{2})=s+t$.

### v2

@@ 1.3. The direct sum of q-matroids
In [10] the notion of direct sum of two $q$-matroids has been introduced. We are going to define the direct sum of $t$ many $q$-matroids iteratively, thanks to the associativity property of the direct sum, proved in [18]. We use Notation 1.7 and the notation used in [17].
@@
Theorem 1.15 ([19, Theorem 3.4]).
Let $\mathcal{M}_{i}=(\mathbb{F}_{q}^{n_{i}},\rho_{i}),i=1,2,$ be $q$-matroids of rank $k_{i}$, and let $\mathcal{M}=\mathcal{M}_{1}\oplus\mathcal{M}_{2}$. Suppose $\mathcal{M}$ is representable over $\mathbb{F}_{q^{m}}$. Then $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are representable over $\mathbb{F}_{q^{m}}$ and $\mathcal{M}=(\mathcal{S}_{1}\oplus\mathcal{S}_{2},\rho_{\mathcal{S}_{1}\oplus\mathcal{S}_{2}})$ for some $n_{i},k_{i}]_{q^{m}/q}$ systems $\mathcal{S}_{i}$, $i\in\{1,2\}$. Furthermore, $(\mathcal{S}_{i},\rho_{\mathcal{S}_{i}})$ is an $\mathbb{F}_{q^{m}}$-representation of $\mathcal{M}_{i}$ for each $i\in\{1,2\}$.
@@
Example 1.17 ([19, Proposition 3.7]).
Let $\mathbb{F}_{q}=\mathbb{F}_{2}$ and $\mathbb{F}_{4}=\{0,1,\omega,\omega+1\}$. Consider the full $4,2]_{4/2}$ system $\mathcal{S}=\mathbb{F}_{4}^{2}$ and set $\mathcal{M}:=(\mathcal{S},\rho_{\mathcal{S}})$. Then $\mathcal{M}\oplus\mathcal{M}$ is not representable.
@@
Let $\mathcal{M}=(\mathbb{F}_{q}^{n},\rho)$ be a $q$-matroid of rank $k$, with $0<k<n$. Then $\mathcal{M}=\mathcal{U}_{k,n}(q)$ if and only if $\langle 0\rangle$ and $\mathbb{F}_{q}^{n}$ are the only cyclic flats of $\mathcal{M}$.
@@ 2. The direct sum of uniform q-matroids
In this section we take the first steps toward the geometric characterization of the representability of the direct sum of $t$ uniform $q$-matroids $\mathcal{U}_{k_{1},n_{1}}(q)\oplus\cdots\oplus\mathcal{U}_{k_{t},n_{t}}(q)$, by characterising its independent spaces. In the rest of the paper we will assume that none of the $t$ summands is the trivial or the free uniform $q$-matroid. This is due to the following observation.
@@
Furthermore, in [19, Proposition 3.9] it is proven that a $q$-matroid $\mathcal{M}$ is $\mathbb{F}_{q^{m}}$-representable as $\mathcal{M}=(\mathcal{S},\rho_{\mathcal{S}})$, if and only if $\mathcal{M}\oplus\mathcal{U}_{s,s}(q)$ is $\mathbb{F}_{q^{m}}$-represented by $(\mathcal{S}\oplus\mathbb{F}_{q}^{s},\rho_{\mathcal{S}\oplus\mathbb{F}_{q}^{s}})$.
@@
For the rest of the paper, assume $1\leq k_{i}<n_{i}$ and let $\mathcal{U}_{k_{i},n_{i}}(q)$, $i\in[t]$ be uniform $q$-matroids.
@@
Let $k_{1},\ldots,k_{t},n_{1},\ldots,n_{t}$ be positive integers, with $1\leq k_{i}<n_{i}$ for $i\in[t]$. Let $\mathcal{M}=\mathcal{U}_{k_{1},n_{1}}(q)\oplus\cdots\oplus\mathcal{U}_{k_{t},n_{t}}(q)$ and for each $\mathcal{J}\subseteq[t]$ denote $k_{\mathcal{J}}:=\sum_{j\in\mathcal{J}}k_{j}$. Then $I\in\mathcal{I}(\mathcal{M})$ if and only if
@@
Let $\mathbf{k}=(k_{1},\ldots,k_{t})\in\mathbb{N}^{t}$ and let $h,k$ be positive integers such that $k=k_{1}+\ldots+k_{t}$ and $1\leq h\leq k-1$, then we define
[…]
@@
Let $k_{1},\ldots,k_{t},k,n_{1},\ldots,n_{t},m,r$ be positive integers, with $1\leq k_{i}<n_{i}\leq m$ for $i\in[t]$ and $k=k_{1}+\ldots+k_{t}$. Let $\mathcal{S}_{i}$ be a $n_{i},k_{i}]_{q^{m}/q}$ system for $i\in[t]$. If $\mathcal{S}=\mathcal{S}_{1}\oplus\ldots\oplus\mathcal{S}_{t}$ is $(\Lambda_{k-1,\mathbf{k}},r)$-evasive, then $\mathcal{S}$ is $(\Lambda_{l,\mathbf{k}},r-k+1+l)$-evasive for any $l\in[k-1]$. In particular, if $\mathcal{S}$ is $(\Lambda_{k-1,\mathbf{k}},k-1)$-evasive, then $\mathcal{S}$ is $(\Lambda_{l,\mathbf{k}},l)$-evasive for any $l\in[k-1]$.
@@
We prove the statement by induction on k1lk-1-l.
If k1l=0k-1-l=0, then there is nothing to prove and the assertion is clearly true.
Suppose that l<k1l<k-1 and assume that the claim holds true for $l+1$. Let us take any subspace $T\in\Lambda_{l,\mathbf{k}}$. We will show that there exists a subspace $T^{\prime}\in\Lambda_{l+1,\mathbf{k}}$ containing $T$ and such that
@@
$\dim_{\mathbb{F}_{q}}(T\cap\mathcal{S})\leq\dim_{\mathbb{F}_{q}}(T^{\prime}\cap\mathcal{S})-1\leq r-k+1+l+1-1=r-k+1+l.$
@@
Since $\mathcal{S}_{i}$ is an $n_{i},k_{i}]_{q^{m}/q}$ system contained in $\iota_{i}(\mathbb{F}_{q^{m}}^{{k_{i}}})$ there exists $v\in\mathcal{S}_{i}\setminus T$. Clearly, $T^{\prime}=T+\langle v\rangle_{\mathbb{F}_{q^{m}}}\in\Lambda_{l+1,\mathbf{k}}$ and satisfies (5).
Case II. Now assume that for every $i\in[t]$ we have that
@@
So, $l\geq k-t$ and $T\supseteq(T\cap\iota_{1}(\mathbb{F}_{q^{m}}^{{k_{1}}}))+(T\cap\iota_{2}(\mathbb{F}_{q^{m}}^{{k_{2}}}))+\ldots+(T\cap\iota_{t}(\mathbb{F}_{q^{m}}^{{k_{t}}}))$.
For every $i\in\{2,\ldots,t\}$, consider $T_{i}=\iota_{1}(\mathbb{F}_{q^{m}}^{{k_{1}}})+\ldots+\iota_{i}(\mathbb{F}_{q^{m}}^{{k_{i}}})$. Then, $\dim_{\mathbb{F}_{q^{m}}}(T_{i})=k_{1}+\ldots+k_{i}$ and, since $T\cap T_{i}\supseteq(T\cap T_{i-1})+(T\cap\iota_{i}(\mathbb{F}_{q^{m}}^{k_{i}}))$,
$\dim_{\mathbb{F}_{q^{m}}}(T\cap T_{i})\in\{\dim_{\mathbb{F}_{q^{m}}}(T\cap T_{i-1})+k_{i}-1,\ldots,k_{1}+\ldots+k_{i}-1\}.$ (6)
@@
Case II.a. Assume that there exists $i\in\{1,\ldots,t\}$ such that $\dim_{\mathbb{F}_{q^{m}}}(T\cap T_{i})=\dim_{\mathbb{F}_{q^{m}}}(T\cap T_{i-1})+k_{i}-1$. Hence,
$T\cap T_{i}=(T\cap T_{i-1})+(T\cap\iota_{i}(\mathbb{F}_{q^{m}}^{{k_{i}}})).$
Let $v_{1}\in\mathcal{S}_{1}\oplus\ldots\oplus\mathcal{S}_{i-1}$ and $v_{i}\in\mathcal{S}_{i}$ such that $v_{1},v_{i}\notin T$. Note that such vectors exist, otherwise $T$ would contain at least one of the $\mathcal{S}_{1},\ldots,\mathcal{S}_{i}$, a contradiction as $T\in\Lambda_{l,\mathbf{k}}$. Also, we have that $\langle v_{1},v_{i}\rangle_{\mathbb{F}_{q^{m}}}\cap T=\{0\}$. Indeed, if $\alpha v_{1}+\beta v_{i}\in\langle v_{1},v_{i}\rangle_{\mathbb{F}_{q^{m}}}\cap T$ with $\alpha,\beta\in\mathbb{F}_{q^{m}}$, and $(\alpha,\beta)\neq(0,0)$, then $\alpha v_{1}\in T\cap T_{i-1}$ and $\beta v_{i}\in T\cap\iota_{i}(\mathbb{F}_{q^{m}}^{{k_{i}}})$, a contradiction as $v_{1},v_{i}\notin T$. Therefore, $T^{\prime}=T\oplus\langle v_{1}+v_{i}\rangle_{\mathbb{F}_{q^{m}}}\in\Lambda_{l+1,\mathbf{k}}$ and satisfies (5).
Case II.b. Suppose now that for every $i\in\{2,\ldots,t\}$ we have $\dim_{\mathbb{F}_{q^{m}}}(T\cap T_{i})\geq\dim_{\mathbb{F}_{q^{m}}}(T\cap T_{i-1})+k_{i}$. Observe that this implies that $\dim_{\mathbb{F}_{q^{m}}}(T\cap T_{2})\geq\dim_{\mathbb{F}_{q^{m}}}(T\cap T_{1})+k_{2}=k_{1}+k_{2}-1$, and inductively, for every $i\in[t]$, $\dim_{\mathbb{F}_{q^{m}}}(T\cap T_{i})\geq k_{1}+\ldots+k_{i}-1$, meaning equality due to (6). Hence, the equality also holds for $i=t$, that is,
$\dim_{\mathbb{F}_{q^{m}}}(T\cap T_{t})=k_{1}+\ldots+k_{t}-1=k-1,$
which is a contradiction to the hypothesis that $\dim_{\mathbb{F}_{q^{m}}}(T)=l<k-1$. ∎
@@
Let $k_{1},\ldots,k_{t},k,n_{1},\ldots,n_{t},n,m$ be positive integers, with $1\leq k_{i}<n_{i}\leq m$ for $i\in[t]$ and $k=k_{1}+\ldots+k_{t}$, $n=n_{1}+\ldots+n_{t}$. Let $\mathcal{S}_{i}$ be an $\mathbb{F}_{q^{m}}$-representation for $\mathcal{U}_{k_{i},n_{i}}(q)$ for each $i\in[t]$. Define the $n,k]_{q^{m}/q}$ system
@@
$(1)\Longleftrightarrow(2)$: Let us call $\mathcal{M}_{1}=(\mathcal{S},\rho_{\mathcal{S}})$ and $\mathcal{M}_{2}=\mathcal{U}_{k_{1},n_{1}}(q)\oplus\ldots\oplus\mathcal{U}_{k_{t},n_{t}}(q)$. Then, $\mathcal{M}_{1}\cong\mathcal{M}_{2}$ if and only if there exists an invertible $\mathbb{F}_{q}$-linear map $\psi:\mathbb{F}_{q}^{n_{1}+\ldots+n_{t}}\longrightarrow\mathcal{S}$ such that $\mathcal{I}(\mathcal{M}_{1})=\mathcal{I}(\psi(\mathcal{M}_{2}))$. Let $G$ be any generator matrix associated to the $n,k]_{q^{m}/q}$ system $\mathcal{S}$ of the form
@@
Case I. If $\Gamma$ does not contain any of the $\mathbb{F}_{q^{m}}$-subspaces $\iota_{i}(\mathbb{F}_{q^{m}}^{{k_{i}}})$, then $\ell:=\rho_{\mathcal{S}}(I)<k$ and $\Gamma\in\Lambda_{\ell,\mathbf{k}}$. This implies that $\mathcal{S}$ is not $(\Lambda_{\ell,\mathbf{k}},\ell)$-evasive, contradicting the hypothesis of $\mathcal{S}$ being $(\Lambda_{k-1,\mathbf{k}},k-1)$-evasive, due to Lemma 2.3. Furthermore, this also shows that for every $\mathbb{F}_{q}$-subspace $I^{\prime}\subseteq\mathcal{S}$, if $\Gamma^{\prime}:=\langle I^{\prime}\rangle_{\mathbb{F}_{q^{m}}}$ does not contain any space of the form $\iota_{i}(\mathbb{F}_{q^{m}}^{{k_{i}}})$ for $i\in[t]$, then
$\rho_{\mathcal{S}}(I^{\prime})=\min\{k,\dim_{\mathbb{F}_{q}}(I^{\prime})\},$ (8)
otherwise, as before, we can apply again Lemma 2.3 and get a contradiction. In fact, we just proved that if $\rho_{\mathcal{S}}(I^{\prime})<\dim_{\mathbb{F}_{q}}(I^{\prime})$ then $\rho_{\mathcal{S}}(I^{\prime})=k$.
Case II. Up to a permutation of the set $[t]$, we can assume without loss of generality that $\Gamma$ contains the $\mathbb{F}_{q^{m}}$-subspaces $\iota_{i}(\mathbb{F}_{q^{m}}^{{k_{i}}})$, for each $i\in[a]$, for some $a$ such that $0<a<t$ (note that $a=0$ is Case I, and $a=t$ was already proved before). Let $\dim_{\mathbb{F}_{q}}(I)=r+s+d$ with $r=\dim_{\mathbb{F}_{q}}(I\cap\mathcal{S}_{[a]})\leq k_{[a]}$ and $s=\dim_{\mathbb{F}_{q}}(I\cap\mathcal{S}_{[t]\setminus[a]})\leq k_{[t]\setminus[a]}$. $I$ can be written as
$I=\langle(v_{1}\mid 0),\ldots,(v_{r}\mid 0)\rangle_{\mathbb{F}_{q}}+\langle(0\mid u_{1}),\ldots,(0\mid u_{s})\rangle_{\mathbb{F}_{q}}+\langle(w_{1}\mid z_{1}),\ldots,(w_{d}\mid z_{d})\rangle_{\mathbb{F}_{q}}.$
@@
    $X_{1}:=\langle v_{1},\ldots,v_{r},w_{1},\ldots,w_{d}\rangle_{\mathbb{F}_{q}}\subseteq\mathcal{S}_{[a]}$ and $\dim_{\mathbb{F}_{q}}(X_{1})=r+d$;
@@
    $X_{2}:=\langle u_{1},\ldots,u_{s},z_{1},\ldots,z_{d}\rangle_{\mathbb{F}_{q}}\subseteq\mathcal{S}_{[t]\setminus[a]}$ and $\dim_{\mathbb{F}_{q}}(X_{2})=s+d$.